# boxthm]Proposition boxthm]Example boxthm]Corollary boxthm]Lemma boxthm]Definition boxthm]Assumption

### v1

@@ 4.2 Fair OT via Cost Learning
$\mathbb{E}\big[\big\lvert\int\varphi d(\mathbf{\Pi}_{\epsilon}(c_{\theta})-\pi^{\star}_{\epsilon}(c_{\theta}))\big\rvert\big]\lesssim\frac{\exp\big(5R_{\Theta}\epsilon^{-1}\big)\lVert\varphi\rVert_{L^{\infty}(\mu\otimes\eta)}}{\sqrt{n}}.$
Then, using the identity $\lvert a^{2}-b^{2}\rvert\leq\lvert a+b\rvert\lvert a-b\rvert$, we show that
$\lvert\mathcal{L}_{\mathbf{F}}(\mathbf{\Pi}_{\epsilon}(c_{\theta}))-\mathcal{L}_{\mathbf{F}}(\pi^{\star}_{\epsilon}(c_{\theta}))\rvert\lesssim\sum_{s=1}^{K_{s}}\sum_{w=1}^{K_{w}}\lvert\textnormal{Tr}\big(\mathbf{\Pi}_{\epsilon}(c_{\theta})^{\top}\mathbf{B}_{uv}\big)-\int_{\mathcal{X}\times\{s\}\times\mathcal{Y}\times\{w\}}d\pi^{\star}_{\epsilon}(c_{\theta})(x,u,y,t)\big\lvert$.
For suitable choice of $\varphi_{sw}\in L^{\infty}(\mu\otimes\eta)$, we have for any $s,w\in[K_{s}]\times[K_{w}]$
$\lvert\textnormal{Tr}\big(\mathbf{\Pi}_{\epsilon}(c_{\theta})^{\top}\mathbf{B}_{uv}\big)-\int_{\mathcal{X}\times\{s\}\times\mathcal{Y}\times\{w\}}d\pi^{\star}_{\epsilon}(c_{\theta})(x,u,y,t)\big\lvert=\big\lvert\int\varphi_{sw}d(\mathbf{\Pi}_{\epsilon}(c_{\theta})-\pi^{\star}_{\epsilon}(c_{\theta}))\big\rvert$
Putting all together yields the result. ∎
This bound provides a uniform control over the expected difference between the fairness loss evaluated on the population and its finite-sample approximation, for any cost function $c_{\theta}$ in the parameterized family. The exponential dependence on $\varepsilon^{-1}$ is unavoidable in the worst case; see rigollet2022sample and altschuler2022asymptotics for related discussions.
Importantly, Section˜4.2 implies that, for sufficiently large training and test samples, the transport plan induced by the learned cost achieves a level of fairness on the test samples similar to that observed on the training samples. In other words, the learned cost can be reused to match new samples with fairness guarantees.
@@ 4.3 Comparison of Both Approaches
Convexity.
The penalized OT problem is convex, whereas the cost learning approach involves a bilevel optimization, which is non-convex and may get stuck in local minima.
Flexibility and Interpretability.
Cost learning offers greater flexibility: the baseline cost function and the discrepancy $\mathscr{D}$ can be chosen by the user. Moreover, selecting an appropriate parameterized family can make the learned cost interpretable.
Reusability.
Penalized OT requires solving a new problem for each batch of samples. In contrast, a learned cost can be reused to match new samples via a standard OT problem.
@@ 5 Experiments
5.1 Approaches and Metrics
Approaches.
We compare our two relaxed approaches: penalized OT and cost learning. For cost learning, we evaluate two parameterizations: a Mahalanobis cost (˜4.1) and a neural cost (˜4.2), where $\phi_{\theta}$ is a multilayer perceptron (MLP) or a ResNet trained from scratch.
5.2 Approaches and Metrics
Approaches.
We compare our two relaxed approaches: penalized OT and cost learning. For cost learning, we evaluate two parameterizations: a Mahalanobis cost (˜4.1) and a neural cost (˜4.2), where $\phi_{\theta}$ is a multilayer perceptron (MLP) or a ResNet embedding.
Metrics.
Following standard practice in fairness research, we analyze the utility–fairness trade-off by visualizing performance in the two-dimensional (utility, fairness) plane. The Pareto front corresponds to all points that, for any given level of fairness, achieve the lowest level of bias. The Pareto front consists of points that, for a given level of fairness, achieve the lowest possible bias; points on this front are considered optimal, as improving one metric requires trading off the other. Bias is measured relative to the transport plan obtained by standard (non-fair) OT with the same entropic regularization, using two quantities: (a) the KL divergence between transport plans, and (b) the difference in total transport cost. In all experiments, the target $\mathbf{F}$ has all coefficients equal to zero, except for the anti-diagonal entries, which are set to $\frac{1}{KsKw}$.
Implementation.
Our experiments are conducted using the Python Optimal Transport library [flamary2021pot, flamary2024pot], which provides implementations of the Sinkhorn and generalized conditional gradient algorithms. In the cost-learning approach, we use iterative differentiation by backpropagating through the Sinkhorn iterations [Pauwels2023] to compute gradients of the bilevel objective using the PyTorch framework [Paszke2019].
@@ 5.3 Comparing Exact and Penalized Approaches
To illustrate the need to move from strict group fairness constraints to relaxed ones, we compare the transport cost and the level of fairness achieved for the transport plans obtained using our FairSinkhorn algorithm, introduced in Section˜3.2, the transport plan obtained using the original Sinkhorn algorithm, and the fairness-penalized optimal transport presented in Section˜4.1, for different values of the fairness penalty. In Fig.˜2, one observes that FairSinkhorn achieves perfect fairness, but with high transport cost. In contrast, Sinkhorn achieves low transport cost with high fairness loss. One observes that adopting the relaxed formulation allows to attain low fairness violations while achieving a better transport cost. This demonstrate that the penalization enables to tune the fairness-utility trade-off of interest.
Figure 2: Comparison of FairSinkhorn , penalized optimal transport, and Sinkhorn on the Gaussian data introduced in Section˜5.4. FairSinkhorn achieves perfect fairness with a high transport cost while Sinkhorn achieves low transport cost with low fairness. The penalized OT interpolates between Sinkhorn and FairSinkhorn .
@@ 5.4 Comparing Penalized OT and Cost Learning
Figure 3: Synthetic Gaussian data experiment. Left: 2D gaussians with binary sensitive attribute $S\in\{0,1\}$ and $W\in\{0,1\}$. Right: Utility/fairness tradeoffs. On the top, the utility metric is the KL divergence between the OT plan with and without the fairness constraint in function of the fairness regularization parameter. On the bottom, the difference between the transport cost with fair and unfair transport plan. The penalized OT correspond to our first relaxed approach while de Mahalanobis, MLP and ResNet ones correspond to the cost learning approach.
Figure 4: Transport plan we get using different level of fairness parameter for each approach with the Gaussian data.
We now evaluate both approaches on simulated datasets. We consider two types of data. First, we study mixtures of Gaussians, where each group is assigned to a distinct mixture component (see Fig.˜3, left). Second, we consider a dataset composed of concentric circles (see Fig.˜5, left), which allows us to examine the effect of data geometry. In both settings, the penalized OT approach achieves a better utility–fairness trade-off, benefiting from its convex formulation. Moreover, for the cost-learning approach, we observe that using more expressive cost models enables lower levels of fairness loss to be attained.
Figure 5: Synthetic circular data experiment. Left: 2D gaussians with sensitive attribute $S\in\{0,1,2\}$ and $W\in\{0,1,2\}$. Right: Utility/fairness tradeoffs. On the top, the utility metric is the KL divergence between the OT plan with and without the fairness constraint in function of the fairness regularization parameter. On the bottom, the difference between the transport cost with fair and unfair transport plan. The penalized OT correspond to our first relaxed approach while de Mahalanobis, MLP and ResNet ones correspond to the cost learning approach.
Figure 6: Transport plan we get using different level of fairness parameter for each approach with the circular data.
@@ 5.5 Assessing the Reusability of Learned Costs
One of the peaks of our cost learning driven approach is that once learned, the cost function can be called on new samples without any retraining. In contrast, the penalized optimal transport plan cannot be reused, and we need to solve a new transport problem for every new batch of data points.
We quantify the reusability of our bilevel algorithm in the following way. We first learn a Mahalanobis cost function $\widehat{\mathbf{M}}$ on a set of $n_{0}=200$ datapoints with $k_{s}=k_{w}=2$ distinct classes. We subsequently sample $k=50$ new datapoints from the same distribution, and $a)$ solve the penalized OT problem on this fresh dataset $b)$ solve the optimal transport problem using $\widehat{\mathbf{M}}$ to define the cost and compute the fairness of the resulting transport plan. Results are displayed in Figure 7. One observes that using the learned cost allows new data points to be matched with significantly lower inference time than the penalized approach, since the latter requires solving the entire penalized optimal transport problem each time new samples are added to the dataset.
Figure 7: We assess the recyclability of the learned cost function by $a)$ learning a cost function on $n_{0}=200$ samples and $b)$ computing the entropic transport plan on batches of $n_{k}$ new unseen samples. Concurrently to this setup, we train the penalized transport plan from scratch for every new batch. Here, $\varepsilon=2$, $n_{k}=100$, $d=2$ and $k_{s}=k_{w}=2$. We set the penalty for all three approaches to yield comparable levels of fairness, but emphasize that our goal in this experiment is to compare the running times of the different approaches.
@@ 6 Conclusion
In this paper, we introduced a novel notion of group fairness for optimal transport. Building on this new notion, we proposed an algorithm that solves entropic regularized optimal transport with perfect fairness, though this can lead to solutions with high transport cost. To achieve a better utility–fairness trade-off, we then propose two relaxed approaches: (i) a convex penalization of the entropic optimal transport objective, and (ii) a cost-learning-based method. For the penalized approach, we derive a sample complexity bound. For the cost learning approach, we bound the expected deviation of the fairness loss when the transport plan is estimated from a finite number of samples drawn from the underlying distributions. Deriving a generalization bound for the cost learning approach is left for future work.
@@ Acknowledgments.
Appendix A A Modified Sinkhorn Algorithm for Fair Optimal Transport
@@
A fair optimal transport plan can hence be obtained using a modified Sinkhorn algorithm which we call FairSinkhorn. Define the map
$\Phi(\mathbf{u},\mathbf{v}):=\bigg(\textnormal{Tr}\Big[\Big(\big[\mathbf{u}_{i}\mathbf{v}_{j}\big]_{ij}\odot\mathbf{B}_{sw}\Big)\big(\mathbf{K}\odot\mathbf{B}_{sw}\big)^{\top}\Big]\bigg)_{sw}.$
The algorithm FairSinkhorn combines projections on the set of marginal constraints with a projection on the set of fairness constraints, highlighted in blue.
Remark 1.
Since FairSinkhorn alternates between projections on sets defined by linear constraints, it is guaranteed to converge [benamou2015iterative].
Appendix B Proofs
B.1 Proof of Section˜3.2
Lemma B.1.
@@ Appendix B Proofs
To finish, just note that $\Pi_{\text{fair}}^{\mathbf{F}}=\Pi(\mu,\eta)\cap\Lambda$ is the intersection of two compact sets (with respect to weak topology) and Lemma B.1 establishes that it is non-empty. Continuity of the map defining the transport problem is enough to conclude the existence of a minimizer. Uniqueness is then a consequence of the strict convexity of KL. ∎
B.2 Proof of Section˜4.1
@@
To finish the lower bound, note that $\mathbb{E}[|Y_{n}-m^{\star}_{\infty}|]\leq\mathcal{O}(\sqrt{n}^{-1})$ follows from (15) provided we show that $m_{\infty}^{\star}=\widehat{m}_{\infty}^{\star}-\lambda\mathcal{L}_{\mathbf{F}}(\pi_{\infty}^{\star})$; this is the content of Lemma B.2 that can be found in Appendix B.2.1.
@@
Lemma B.2 shows that $m^{\star}_{\infty}=\widehat{m}_{\infty}^{\star}-\lambda\mathcal{L}_{F}(\pi^{\star}_{\infty}\rangle$ and, similar to the lower bound part of the proof (see (15)), the sample complexity of entropic optimal transport implies that,
@@
To obtain these two inequalities and complete the proof, it suffices to note that the conclusion of Lemma B.2 allows us to apply Lemma˜B.4 (sub-Gaussian concentration, which is stronger, readily implies the expectation bound used here), which yields the desired inequality (17). To obtain inequality (18) just note that, for $s,w\in[K_{s}]\times[K_{w}]$,
@@
and the bound in expectation is once more a matter of invoking Lemma˜B.4. ∎
B.2.1 Technical Lemmas
Lemma B.2.

### v2

@@ 4.2 Fair OT via Cost Learning
$\mathbb{E}\big[\big\lvert\textstyle\int\varphi d(\mathbf{\Pi}_{\epsilon}(c_{\theta})-\pi^{\star}_{\epsilon}(c_{\theta}))\big\rvert\big]\lesssim\frac{\exp(5R_{\Theta}\epsilon^{-1})\lVert\varphi\rVert_{L^{\infty}(\mu\otimes\eta)}}{\sqrt{n}}.$
Then, using $\lvert a^{2}-b^{2}\rvert\leq\lvert a+b\rvert\lvert a-b\rvert$, we show that
$\lvert\mathcal{L}_{\mathbf{F}}(\mathbf{\Pi}_{\epsilon}(c_{\theta}))\!-\!\mathcal{L}_{\mathbf{F}}(\pi^{\star}_{\epsilon}(c_{\theta}))\rvert\!\lesssim\!\sum_{s=1}^{K_{s}}\sum_{w=1}^{K_{w}}\big\lvert\textnormal{Tr}\big(\mathbf{\Pi}_{\epsilon}(c_{\theta})^{\top}\mathbf{B}_{uv}\big)$
$\hskip-160.00024pt-\textstyle\int_{\mathcal{X}\times\{s\}\times\mathcal{Y}\times\{w\}}d\pi^{\star}_{\epsilon}(c_{\theta})(x,u,y,t)\big\lvert$.
Then, for a suitable choice of $\varphi_{sw}\in L^{\infty}(\mu\otimes\eta)$, we have for any $s,w\in[K_{s}]\times[K_{w}]$
$\big\lvert\textnormal{Tr}\big(\mathbf{\Pi}_{\epsilon}(c_{\theta})^{\top}\mathbf{B}_{uv}\big)-\textstyle\int_{\mathcal{X}\times\{s\}\times\mathcal{Y}\times\{w\}}d\pi^{\star}_{\epsilon}(c_{\theta})(x,u,y,t)\big\lvert$
$\hskip-129.00012pt=\big\lvert\textstyle\int\varphi_{sw}d(\mathbf{\Pi}_{\epsilon}(c_{\theta})-\pi^{\star}_{\epsilon}(c_{\theta}))\big\rvert$.
Putting everything together yields the result. ∎
This bound provides a uniform control over the expected difference between the fairness loss evaluated on the population and its finite-sample approximation. The exponential dependence on $\varepsilon^{-1}$ is unavoidable in the worst case; see rigollet2022sample and altschuler2022asymptotics for related discussions. Importantly, Section˜4.2 implies that, for two sufficiently large samples, the transport plan computed with a given cost achieves a comparable level of fairness. In other words, the same cost function can be reused to match new samples with fairness guarantees.
@@ 4.3 Comparison of Both Approaches
Convexity. The penalized OT problem is convex, whereas the cost learning approach involves a bilevel optimization, which is non-convex and may get stuck in local minima.
Flexibility. The cost learning approach restricts the transport map to entropic-regularized solutions with a parametric cost. This problem is thus structurally more constrained than the penalized formulation, which allows solutions outside the set of optimal couplings and is therefore more flexible.
Reusability and interpretability. Penalized OT requires solving a new problem for each batch of samples. In contrast, a learned cost can be reused to match new samples via a standard OT problem. Moreover, when the cost belongs to an interpretable family, it can yield insight into how specific features contribute to (un)fairness.
@@ 5 Experiments
5.1 Experimental Setup
Data generation. Building on the illustrating example of segregated school systems (Examples 3.13.2), we consider two synthetic problems. For the first one (Fig.˜2, left), the positions of students ($X$) and schools ($Y$) are drawn according to a mixture of two Gaussian distributions such that privileged students ($S=1$) are closer to elite schools ($W=1$), and underprivileged students closer to non-elite schools. For the second one (Fig.˜2, right), privileged students and elite schools are drawn from a centered Gaussian distribution, while underprivileged students and non-elite schools are sampled on a centered circle of radius 2. In both cases, there are 10 times more schools than students, and sensitive attributes are balanced.
Figure 2: Simulated datasets: Gaussians (left) and Circles (right).
Fairness target. We aim to assign approximately $60\%$ of underprivileged students to elite schools, which corresponds to the fairness target $\mathbf{F}=\begin{bmatrix}0.20&0.30\\ 0.28&0.22\end{bmatrix}$.
Approaches. We compare our three approaches: FairSinkhorn , penalized OT and cost learning. For cost learning, we evaluate two parameterizations: a Mahalanobis cost (˜4.1) and a neural cost (˜4.2), where $\phi_{\theta_{1}}$ and $\phi_{\theta_{2}}$ are multilayer perceptrons (MLP). Implementation details are available in Appendix˜D. We also include vanilla entropic OT with varying $\varepsilon$ as a baseline.
Metrics. Following standard practice in fairness research, we analyze the cost–fairness trade-off by visualizing performance in the two-dimensional (cost, fairness) plane. Cost is measured relative to the transport plan obtained by standard (non-fair) OT with the same entropic regularization, using the difference in total transport cost.
Figure 3: a. FairSinkhorn achieves perfect fairness with a high transport cost while Sinkhorn achieves low transport cost with low fairness. The penalized OT interpolates between Sinkhorn and FairSinkhorn . b. Cost-fairness trade-off of our penalized and cost-learning approaches on both datasets for varying fairness penalties. c. Vanilla entropic OT with different values of $\varepsilon$ is included as a baseline. We also display the fairness targets reached at a few selected points.
@@ 5.2 Comparing Exact and Penalized Approaches
To illustrate the need to move from exact group fairness constraints to relaxed ones, we compare, on the Gaussian dataset, the (cost,fairness) trade-off achieved by the original Sinkhorn algorithm, our FairSinkhorn algorithm (Section˜3.2), and our fairness-penalized approach (Section˜4.1) for varying values of the fairness penalty. As shown in Fig.˜3 a., our penalized approach smoothly interpolates between Sinkhorn , which attains low transport cost but large fairness violations, and FairSinkhorn , which enforces perfect fairness at the expense of higher transport cost. This relaxed formulation thus enables fine-grained control over the fairness-cost trade-off.
@@ 5.3 Comparing Penalized OT and Cost Learning
Fig.˜3 b. compares our two relaxed approaches across our two datasets, allowing us to highlight the effect of data geometry. For the Gaussian dataset, the cost-fairness trade-off of penalized OT, cost learning with a Mahalanobis metric, and cost learning with an MLP are similar. On the Circles dataset, however, the penalized approach achieves a superior trade-off, reaching any level of fairness at lower cost than the cost learning methods. This is expected, as penalized OT imposes no restrictions on the transport plans (see Section 4.3). Moreover, cost learning with the Mahalanobis parameterization cannot reduce fairness loss below $10^{-2}$, whereas the MLP variant is flexible enough to transform the data nonlinearly, which is required to closely match the fairness target on this problem. Finally, for both datasets, increasing the entropic penalty $\varepsilon$ in vanilla OT fails to achieve the desired fairness, highlighting the relevance and effectiveness of our approaches.
Figure 4: Left: Inference time of penalized OT and cost-learning approaches, highlighting the much faster inference of the cost-learning methods once the cost function is learned. Right: Fairness levels achieved on new samples using the learned cost function, with vanilla OT shown as a baseline. Leveraging the learned cost allows new samples to attain fairness comparable to that observed during training.
@@ 5.4 Assessing the Reusability of Learned Costs
A key advantage of our cost learning approach is that, once the cost function is learned, it can be applied to match new samples using vanilla OT without retraining. In contrast, the penalized optimal transport plan cannot be reused, requiring a new penalized OT problem to be solved for each incoming batch of data. In this section, we quantify the benefits of reusing the learned cost on new samples in terms of inference time and fairness generalization. To this end, we first learn a Mahalanobis cost and an MLP-based cost on the Gaussian problem, using a training dataset of $1000$ samples from $X$ and $100$ samples from $Y$. We subsequently draw test datasets of $500$ new samples from $X$ and $50$ new samples from $Y$, solve the vanilla entropic OT problem using the learned costs, and record both the inference time and fairness loss of the resulting transport plan for each test dataset. For comparison, we also report the inference time of penalized OT and the fairness of vanilla OT on the same test data. Results are shown in Figure 4. We observe that using the learned cost allows matching new samples with significantly lower inference time than the penalized approach since the latter requires solving the full penalized optimal transport problem each time new samples are added to the dataset. Furthermore, fairness levels on new samples remain close to those achieved during training. While a generalization gap is present—larger for the MLP cost than for the Mahalanobis cost, as expected—this demonstrates that the learned costs generalize to new samples and achieve substantially better fairness than vanilla OT.
@@ 6 Conclusion
We introduced a novel notion of group fairness for optimal transport and propose an algorithm that computes entropic regularized transport with perfect fairness. To improve the cost–fairness trade-off, we developed two relaxed approaches: (i) a convex penalization of the entropic OT objective, and (ii) a cost-learning-based method. Both approaches are supported by theoretical results. We hope our work will foster further research on fairness constraints in optimal transport. A promising direction for future work is to study how fairness constraints in OT can lead to fairer outcomes downstream.
@@ Acknowledgments.
Appendix A Extended Related Work
A.1 Fairness in Matching Mechanisms
Since the pioneering work of gale1962college on the stable matching problem, a series of works have provided theoretical studies of the fairness in matching mechanisms. A first line of research focuses on the stability of GS-like algorithms under constraints akin to individual fairness, ensuring that similar individuals receive similar outcomes by the decision-making process. For instance, karni2021fairness analyze the Gale-Shapley (GS) algorithm through the lens of preference informed individual fairness. This criterion allows for small deviations from individual fairness, which requires that similar individuals should be treated similarly. They devise a fair variant of the GS algorithm that ensure both approximately fair and stable matchings. Similarly, devic2023fairness analyze merit-based fairness of one-to-one matching mechanisms, a notion close to individual fairness, under partial uncertainty over individual’s merit. A second line of research introduces metrics of fairness closer to group fairness, with a specific focus on kidney paired donation programs, a singular matching problem subject to domain specific constraints — in this setting, donors and patients must be matched by forming cycles due to legal constraints, which leads to a reformulation of the matching problem as a constrained integer programming problem. In their seminal work, ashlagi2014free unveil an unfair market collapse phenomenon in kidney pair donation programs: instead of introducing all patients to the matching system, hospitals internally operate easy matches i.e. do not participate fully in the exchange mechanism. This leads to a loss in transplantation opportunities for highly-sensitized patients that are hard to match. dickerson2014price and following work by zhang2025fairness also consider the case of kidney paired donations programs. Their definition of fairness enforces a minimal amount of program participation for the group of highly-sensitized patients. More recently, lobo2025fair consider a combinatorial multimatching problem under uncertainty: in their setup, the goal is to match agents and resources without exact knowledge of their pairwise valuations. Their definition of fairness relies on the average utility within predefined groups, which is closest to our definition among previous work.
Our contribution differentiates from these works in two aspects. First, we do not consider iterative matching mechanisms that operate at the individual level such as the GS algorithm: our work focuses on matching through (entropic) optimal transport. Secondly, our definition of fairness is different from previously introduced definitions, and highly modular to accommodate fairness preferences of a central planner. Instead of requiring individual fairness, building on individual rankings over possible assignements or participation constraints, we consider a matching problem between two populations (f.i. students and universities) further structured by a sensitive attribute (f.i. privileged and unprivileged students, and elite and non-elite universities). We enforce matching quotas given by a central planner, that specify for each subgroup the proportion of individuals that should be matched to subgroups in the second population.
A.2 Fairness in Bipartite Graph Matching
A closely related topic to our working is bipartite matching in graphs [godsil1981matchings, zdeborova2006number, noiry2021online]. Given a graph $\big(\mathcal{U}\cup\mathcal{V},\mathcal{E}\big)$ made of two sets of vertices $\mathcal{U}$ and $\mathcal{V}$ and edges $\mathcal{E}=(e_{ij})_{i\in\mathcal{U},j\in\mathcal{V}}$, the goal is to select a one-to-one matching between vertices in $\mathcal{U}$ and vertices in $\mathcal{V}$, under the constraint that two vertices can only be matched if they are linked by an edge in $\mathcal{E}$. Some works have considered fairness constraints in this setup. For instance, castera2024price consider this problem with ad-hoc partitions of one side of the graph in disjoint groups, and study the maximal number of individuals that may be matched per group. This is similar to definitions of fairness through participation ratios discussed earlier in the work of zhang2025fairness. Similarly, panda2024individual study a doubly constrained setup in which the set of items $\mathcal{U}$ is structured through groups. They enforce both individual-fairness and group-fairness inspired constraints. In their work, $(a)$ individuals in $\mathcal{U}$ have preferences over individuals in $\mathcal{V}$, and $b)$ every individual in $\mathcal{V}$ can only accept an certain number of individuals from each subgroup in $\mathcal{U}$.
Bipartite matching is hence fundamentally different from our setting and somewhat closer to the literature on fair division [steinhaus1949division, weller1985fair, budish2011combinatorial] since $a)$ we allow for mass splitting between individuals on both sides ; $b)$ individuals are characterized by features in our setting, and not by their relative position within a graph $\mathcal{E}$ ; $c)$ our problem is symmetric in the sens that the two matched measures play the same role.
A.3 Fairness and Optimal Transport
Closer in spirit to our work is the recent article by nguyen2024towards, which imposes fairness constraints directly on a Wasserstein barycenter by requiring approximate equality of sliced-Wasserstein distances SW [nadjahi2020statistical] to multiple marginals $\mu_{1},\dots,\mu_{K}$.
where $\textnormal{{SW}}(\cdot,\cdot)$ is the sliced-Wasserstein distance between distributions [nadjahi2020statistical] and $\varepsilon>0$ is a tolerance threshold. Similarly to their work, we enforce additional constraints on the optimal transport to match it to a definition of fairness based on subgroups. However, we study the problem of computing transport plans that satisfy explicit mass constraints between groups defined by sensitive attributes, a setting that, to the best of our knowledge, has not been previously considered.
A.4 Constrained and Statistical Optimal Transport
Appendix B A Modified Sinkhorn Algorithm for Fair Optimal Transport
@@
Appendix C Proofs
C.1 Proof of Section˜3.2
Lemma C.1.
@@ Appendix C Proofs
To finish, just note that $\Pi_{\text{fair}}^{\mathbf{F}}=\Pi(\mu,\eta)\cap\Lambda$ is the intersection of two compact sets (with respect to weak topology) and Lemma C.1 establishes that it is non-empty. Continuity of the map defining the transport problem is enough to conclude the existence of a minimizer. Uniqueness is then a consequence of the strict convexity of KL. ∎
C.2 Proof of Section˜4.1
@@
To finish the lower bound, note that $\mathbb{E}[|Y_{n}-m^{\star}_{\infty}|]\leq\mathcal{O}(\sqrt{n}^{-1})$ follows from (15) provided we show that $m_{\infty}^{\star}=\widehat{m}_{\infty}^{\star}-\lambda\mathcal{L}_{\mathbf{F}}(\pi_{\infty}^{\star})$; this is the content of Lemma C.2 that can be found in Appendix C.2.1.
@@
Lemma C.2 shows that $m^{\star}_{\infty}=\widehat{m}_{\infty}^{\star}-\lambda\mathcal{L}_{F}(\pi^{\star}_{\infty}\rangle$ and, similar to the lower bound part of the proof (see (15)), the sample complexity of entropic optimal transport implies that,
@@
To obtain these two inequalities and complete the proof, it suffices to note that the conclusion of Lemma C.2 allows us to apply Lemma˜C.4 (sub-Gaussian concentration, which is stronger, readily implies the expectation bound used here), which yields the desired inequality (17). To obtain inequality (18) just note that, for $s,w\in[K_{s}]\times[K_{w}]$,
@@
and the bound in expectation is once more a matter of invoking Lemma˜C.4. ∎
C.2.1 Technical Lemmas
Lemma C.2.